OpenStudy (anonymous):

find the missing probability in the table below : x) P(X) 1) 0.30 2) 0.40 3) 0.15 4) ?

OpenStudy (aum):

There must be more information to this question. Something is missing.

OpenStudy (anonymous):

NO ITS JUST THIS @aum

OpenStudy (here_to_help15):

Let me try:)

OpenStudy (aum):

Can you post a screenshot of the question?

OpenStudy (anonymous):

OK HOLD ON @aum

OpenStudy (anonymous):

@aum

OpenStudy (aum):

There must be something above, maybe a diagram or some other earlier data to which this question applies.

OpenStudy (anonymous):

no there is nothing more than this @aum

OpenStudy (aum):

What was the question previous to this?

OpenStudy (anonymous):

just to find the probability of the chart ! @aum

OpenStudy (aum):

So they are referring to some chart. If you keep going back somewhere you may find something that says to the effect: for the following questions use this chart or diagram or data.

OpenStudy (anonymous):

This was the only chart there @aum

OpenStudy (aum):

I think there is missing info and I am not able to answer with what is provided.

OpenStudy (anonymous):

its ok @aum thanks anyway

OpenStudy (here_to_help15):

I'll try lemme see.:)

OpenStudy (here_to_help15):

All probabilities in a probability distribution must add to 1 So let x = missing probability This then means that 0.3 + 0.4 + 0.15 + x = 1 0.85 + x = 1 x = 1 - 0.85 x = 0.15 So the missing probability is 0.15 Now we're tasked to find P(X > 2) P(X > 2) = P(X = 3) + P(X = 4) P(X > 2) = 0.15 + 0.15

OpenStudy (here_to_help15):

P(X > 2) = 0.3

OpenStudy (here_to_help15):

hope that helped a lil:)

OpenStudy (anonymous):

yes it did but what was the answer @Here_to_Help15

OpenStudy (here_to_help15):

P(X>2)=0.3

OpenStudy (anonymous):

thanks @Here_to_Help15

OpenStudy (here_to_help15):

your welcome :)

OpenStudy (aum):

Shouldn't there be a statement somewhere that states the following table represents a COMPLETE probability distribution?

OpenStudy (here_to_help15):

That is true @aum but all i did was pull what ever i could from this question.

OpenStudy (aum):

You did a good job!

OpenStudy (here_to_help15):

Thank you:)

OpenStudy (anonymous):

thanks to both @aum @Here_to_Help15

OpenStudy (aum):

The problem should have stated somewhere that X can have only 4 values: 1, 2, 3 and 4. That would make it a COMPLETE probability distribution and we can conclude the probabilities will all add up to 1.